OpenStudy (anonymous):

how do you isolate X or Y? (3/x) + (2/y) = 16 (2/x) - (1/y) = -1 I can't seem to succeed in doing it :(

OpenStudy (anonymous):

\[put \frac{ 1 }{ x }=u,\frac{ 1 }{ y }=v and then solve\]

OpenStudy (anonymous):

what do you mean? I only get this: \[3y + 2x = 16xy\]

OpenStudy (anonymous):

3u+2v=16 2u-v=-1 solve for u and v

OpenStudy (anonymous):

finally u=1/x orx=1/u,y=1/v

OpenStudy (loser66):

@surjithayer apply cal3 ?? change variable ?

OpenStudy (anonymous):

3u+2v=16 ...(1) 2u-v=-1 ...(2) (1)+2(2) 3u+2v+4u-2v=-2 7u=-2 u=-2/7 from(2) 2*-2/7-v=-1 -4/7+1=v v=3/7 u=-2/7,1/x=-2/7,x=-7/2 v=3/7,1/y=3/7,y=7/3